OpenStudy (anonymous):

a hat contains 4 marbles one red one green one blue and one yellow. two marbles are sletected at ramdon without replacment. what is the probablity that one is not yellow

OpenStudy (anonymous):

Let's say that during the first drawing you do NOT pull out a yellow. That probablity is 3/4 then that marble is tossed out. Let's say that during the second drawing you DO pull out yellow.. That probablity is 1/3 because there is one yellow and only three marbles left. so 3/4 * 1/3 = 1/4 I think this is correct.

OpenStudy (anonymous):

Thanks for the reply

OpenStudy (anonymous):

The probability that one is not yellow = 1 - probability that one is yellow To calculate the probability that one is yellow = either the first one is yellow or the second one is yellow = (1/4 * 1) + (3/4 * 1/3) = 1/4 +3/12 = 1/2. The probability that neither is yellow = 1-1/2 = 1/2